# Towards Context-aware Reasoning-enhanced Generative Searching in E-commerce

### v2

@@ 4.2. Main Results
Table 2 presents the overall performance and relative improvements of our framework compared with a wide range of baselines. It can be observed that the generative approaches consistently outperform the matching-based baselines across all datasets and metrics. We believe this advantage stems from the autoregressive modeling paradigm of generative models, which captures richer query–item interactions than simple embedding similarity comparison. Moreover, introducing SIDs enables structurally related items to share SID prefixes, preserving semantic proximity and further boosting their performance. In addition, enhanced with complex and rich context as input, the LatentR3 model achieves the strongest results among baselines, underscoring the importance of explicit context modeling in search-based recommendation.
@@
More importantly, our proposed CRS framework consistently achieves superior performance on all three datasets. These results highlight the effectiveness of CRS in deeply understanding user search contexts and inferring user intentions through explicit reasoning. In particular, compared with the best-performing baseline, CRS yields relative improvements of 3–11% on the large-scale datasets (All-100K and All-50K) and up to 47% on the challenging Fashion-27K dataset. Unlike the other two datasets, Fashion-27K contains highly similar items within a single category, making them harder to distinguish by simple query–item correlations. The substantial gains on this dataset further validate the strength of our approach, especially in scenarios where user intent needs to be inferred with fine-grained reasoning.
4.3. Investigation of Self-evolving Post-training
To assess the effectiveness of the proposed self-evolving post-training paradigm, we plot in Figure 2 the model’s performance at each iteration, measured by NDCG@10 and HR@10 on the All-100K and All-50K datasets.
The results clearly show that iterative training substantially boosts the performance of the entire framework. In particular, the relative improvement of NDCG@10 over the distilled model $\mathcal{M}_{\text{sft}}^{0}$ reaches 39.02% and 21.21% on the two datasets, respectively. These gains validate our hypothesis that a simple distillation approach may struggle to generalize due to data–prior and model–capacity biases, whereas the proposed self-evolving paradigm progressively strengthens the model’s ability to perform deep reasoning over complex search contexts.
We also observe that RL updates in some iterations may lead to performance degradation. To investigate this, we compare CRS with a variant trained solely by iterative SFT. This variant consistently yields inferior results, highlighting the necessity of incorporating RL, which enlarges the policy space through deeper exploration and ultimately leads to stronger performance.
$4.$ Ablation Study on R-GRPO
@@ 4.3. Investigation of Self-evolving Post-training
Beyond the self-evolving post-training paradigm, the proposed debiased reinforcement learning algorithm (R-GRPO) also plays a critical role in enhancing the context-aware reasoning ability of the CRS framework. To validate its effectiveness, we conduct an ablation study on the All-50K dataset, comparing the performance of models trained with R-GRPO against those trained with the original GRPO. The results, shown in Table 3, indicate that R-GRPO consistently outperforms GRPO across all three RL training iterations. These findings confirm our expectation and provide strong evidence that R-GRPO effectively mitigates the inherent drawbacks in conventional RL algorithms for search-based recommendation.
@@ 4.5. Generalizability on Insufficient Context
In this section, we investigate the generalizability of the proposed CRS framework under insufficient user context. Specifically, users are grouped into three categories according to the number of their historical queries, representing varying degrees of contextual sufficiency. We then compare the performance of CRS with BGE, DSI, and LatentR3 to examine how explicit context-aware reasoning benefits users with limited interaction history. The results are illustrated in Figure 3.
From the figure, we can first observe that the matching-based BGE model achieves similar performance across all user groups, reflecting its limited sensitivity to contextual richness and indicating the semantic uniformity of the queries. Besides, LatentR3 consistently outperforms DSI, highlighting the necessity of incorporating contextual signals in e-commerce search scenarios. More importantly, the proposed CRS framework delivers substantial gains over baselines models in all user groups. Notably, the results suggest that it excels at boosting the recommendation performance for cold-start users, achieving a relative performance boost of 8.55%. This result suggests that the explicit reasoning process in CRS enables the model to better infer user intent and uncover latent preferences even from insufficient contextual information.
@@
4.6. Scalability Analysis
To further evaluate the scalability of the proposed CRS framework, we scale the backbone model to larger sizes, including 1.7B and 4B parameters, and report the results on the Fashion-27K dataset in Table 4. As the reasoning capacity of the backbone increases with model size, the overall performance of CRS continues to improve across all evaluation metrics, effectively validating the scalability of the proposed framework.

### v3

@@ 4.2. Main Results
Table 2 summarizes the overall performance and relative improvements of our framework compared with a broad set of baselines. As shown, generative approaches consistently outperform matching-based methods across all datasets and evaluation metrics. We attribute this advantage to the autoregressive modeling paradigm of generative models, which captures richer query–item interactions than simple embedding similarity. Furthermore, the introduction of structured item identifiers (SIDs) allows semantically related items to share common SID prefixes, effectively preserving structural and semantic proximity and leading to additional performance gains. Among the baselines, LatentR3 achieves the strongest performance by incorporating rich contextual information as input, highlighting the importance of explicit context modeling in search-based recommendation.
@@
More importantly, our proposed CRS framework consistently delivers superior performance across all three datasets. These results demonstrate the effectiveness of CRS in deeply understanding user search contexts and inferring user intent through explicit reasoning. In particular, compared with the strongest baseline, CRS achieves relative improvements of 3–11% on the large-scale datasets (All-100K and All-50K) and up to 47% on the more challenging Fashion-27K dataset. Unlike the other two datasets, Fashion-27K contains a large number of highly similar items within the same category, making them difficult to distinguish based on simple query–item correlations. The substantial gains observed on this dataset further validate the strength of our approach, especially in scenarios that require fine-grained intent understanding and reasoning.
4.3. Ablation Study on R-GRPO
@@
Beyond the self-evolving post-training paradigm, the proposed debiased reinforcement learning algorithm (R-GRPO) also plays a critical role in enhancing the context-aware reasoning ability of the CRS framework. To validate its effectiveness, we conduct an ablation study on the All-50K dataset, comparing the performance of models trained with R-GRPO against those trained with the baselineGRPO. The results, shown in Table 3, indicate that R-GRPO consistently outperforms GRPO across all three RL training iterations. These findings confirm our expectation and provide strong evidence that R-GRPO effectively mitigates the inherent drawbacks in conventional RL algorithms for search-based recommendation.
@@ 4.4. Generalizability on Insufficient Context
In this section, we further investigate the generalization capability of the proposed CRS framework under scenarios with insufficient user context. Specifically, users in the All-100K dataset are partitioned into three groups based on the number of their historical queries, representing varying levels of contextual availability ranging from cold-start to context-rich users. We compare CRS with representative matching-based and generative baselines, including BGE, DSI, and LatentR3, to examine how explicit context-aware reasoning benefits users with limited interaction history. The results are illustrated in Figure 2.
As shown in the figure, the matching-based BGE model exhibits relatively stable performance across all user groups, suggesting that it is largely insensitive to the amount of available user context and primarily relies on query–item semantic similarity. This behavior also reflects the semantic homogeneity of user queries in the dataset. In contrast, LatentR3 consistently outperforms DSI across all groups, highlighting the importance of incorporating richer contextual signals for modeling user intent in e-commerce search scenarios.
Furthermore, the proposed CRS framework consistently achieves substantial improvements over all baseline models for users at different context levels. Notably, CRS demonstrates a pronounced advantage in cold-start settings, achieving a relative performance gain of 8.55% for users with minimal historical interactions. This observation indicates that the explicit reasoning mechanism in CRS enables more effective inference of user intent and latent preferences by leveraging structured reasoning over limited contextual evidence, rather than relying solely on surface-level query–item correlations. Overall, these results validate the robustness and practical applicability of CRS, particularly in real-world scenarios where user context is often sparse or incomplete.
4.5. Scalability Analysis
To further evaluate the scalability of the proposed CRS framework, we increase the backbone model size to larger configurations with 1.7B and 4B parameters and report the corresponding results on the Fashion-27K dataset in Table 4. As the model capacity grows, the reasoning capability of the backbone is correspondingly enhanced, leading to consistent performance improvements across all evaluation metrics. These results demonstrate that CRS effectively benefits from larger backbone models and validate the strong scalability of the proposed framework.
@@ 5. Conclusion
Acknowledgements.
This research was supported by grants from the National Natural Science Foundation of China (No. 62502486, No. U23A20319), the grants of Provincial Natural Science Foundation of Anhui Province (No. 2408085QF193). This research was also funded by Kuaishou Research Project.
Appendix A Algorithm of R-GRPO
To better clarify the procedure of the training process, we summarize the complete training workflow of R-GRPO in Algorithm 1. Compared to the original GRPO algorithm, our proposal mainly modifies the rollout process and the corresponding reward estimation schema. By explicitly introducing the ranking information, the R-GRPO algorithm focuses on modeling the potential ranking ability of each reasoning trajectory and therefore leads to performance gain for search-based recommendation. Moreover, since the algorithm is independent from the policy optimization procedure, it can be easily extended to diverse variants of recent RL algorithms.
Algorithm 1 R-GRPO: Debiased Reinforcement Learning for Generative Search
1:Training query set $\mathcal{Q}$;
2:Policy model $M_{\theta}$; reference model $M_{\text{ref}}$;
3:Number of reasoning rollouts $G$; beam width $K$;
4:Clipping parameter $\epsilon$; KL coefficient $\beta$
5:for all $q\in\mathcal{Q}$ do
6:  // Stage 1: Reasoning Trajectory Generation
7:  Rollout $G$ reasoning trajectories $\{o_{i}\}_{i=1}^{G}\sim M_{\theta}(\cdot\mid q)$
8:  // Stage 2: Item Beam Search and Reward Assignment
9:  for $i=1$ to $G$ do
10:   Perform beam search conditioned on $(q,o_{i})$
11:   Obtain ranked item list $\{o_{i}^{(j)}\}_{j=1}^{K}$
12:   for $j=1$ to $K$ do
13:     Compute reward $r_{i}^{j}$ using Eq. (3)–(6)
14:   end for
15:   Select optimal rollout $o_{i}^{*}=\arg\max_{j}r_{i}^{(j)}$
16:   Compute rank-aware aggregated reward:
$r_{i}^{*}=\frac{1}{W}\sum_{n=1}^{K}\frac{r_{i}^{n}}{1+\log n},\quad W=\sum_{n=1}^{K}\frac{1}{1+\log n}$
17:  end for
18:  // Policy Optimization
19:  Update policy parameters $\theta$ using standard GRPO by maximizing Eq. (8)
20:end for
Appendix B Investigation of Self-evolving Post-training
To assess the effectiveness of the proposed self-evolving post-training paradigm, we present in Figure 3 the model performance across successive training iterations, evaluated using NDCG@10 and HR@10 on the All-100K and All-50K datasets.
As shown in the figure, iterative post-training substantially and consistently enhances the performance of the overall framework. In particular, compared with the initially distilled model $\mathcal{M}_{\text{sft}}^{0}$, CRS achieves relative improvements of up to 39.02% and 21.21% in terms of NDCG@10 on All-100K and All-50K, respectively. These significant gains strongly validate our hypothesis that a simple distillation approach may struggle to generalize due to data–prior and model–capacity biases. In contrast, the proposed self-evolving paradigm enables the model to progressively refine its reasoning behavior, thereby strengthening its ability to perform deep and context-aware reasoning over complex search scenarios.
We further observe that RL updates in some iterations may introduce temporary performance degradations. To better understand this phenomenon, we compare CRS with a variant trained using only iterative SFT. This variant consistently underperforms the full CRS framework, highlighting the critical role of reinforcement learning, which expands the policy space through exploration and ultimately leads to stronger and more robust performance.
Appendix C Reasoning Trajectories Showcases
We present a comprehensive input-output example with some steps omitted in Table 5 to better illustrate how the CRS framework reason over complex search context and infer the hidden intents of the user. Note that we also observe potential issues of repetition or inconsistency in some reasoning trajectories, which is one of the limitation of the current design.
Table 5. Case study of an input-output example.
Instruction:
The following provides a brief user profile and historical search records of an e-commerce user. Please analyze the user’s click preferences when searching for items from both the user profile analysis and historical search record analysis perspectives, and based on this, predict the SID of the target product.
Input:
 ”user profile”:{”age”:”31-40”, ”resident area”:”northwest, Xining”, …},
 ”history”:[
  {”text”:”iphonewatch strap”, ”time”:”06/10 13:04:30”, ”location”:”Xining”, ”clicked_items”:[], ”non_clicked_items”:
   [{”title”:”Compatible with Apple Watch Series 10/9/8/7 Braided Magnetic Buckle Ultra2 Premium-feel SE Stylish Watch Band”,
   ”brand”:”none”, ”category”:”Smart Device Accessories”, ”price”:”28.8”, ”gmv”:”58.2”, ”SID”:”<a_253><b_490><c_253><d_0>”},
  {”text”:”badminton bag”, ”time”:”06/15 18:28:58”, ”location”:”Xining”, ”clicked_items”:
   [{”title”:”Yonex Badminton Backpack Dual-Shoulder Professional Portable Women’s Model YY 3-Piece Set Large
   Capacity Backpack Men’s Model Shoe Compartment”, ”brand”:”YONEX”, ”category”:”Badminton Bag”, ”price”:”189.0”,
   ”gmv”:”3041.0”, ”SID”:”<a_156><b_105><c_95><d_0>”},
  ”non_clicked_items”:
   [{”title”:”Waterproof Badminton Racket Bag New Large Capacity Sports Backpack Unisex College Student Backpack
   Dual-Shoulder Bag”, ”brand”:”none”, ”category”:”Badminton Bag”, ”price”:”39.99”, ”gmv”:”604.92”,
   ”SID”:”<a_158><b_98><c_139><d_0>”},
  {”text”:”Badminton uniform”, ”time”:”06/15 18:36:34”, ”location”:”Xining”, ”clicked_items”:
   [{”title”:”2024 New Yonex Badminton Apparel Blue and White Porcelain YY Tournament Uniform Quick-Dry Breathable Sports
   Short-Sleeve for Men and Women”, ”brand”:”YONEX”, ”category”:”Badminton Uniform”, ”price”:”35.0”, ”gmv”:”1400.0”,
   ”SID”:”<a_156><b_95><c_95><d_0>”},
  ”non_clicked_items”:
   [{”title”:”Li Zijia Competition-Grade Short-Sleeve Set Badminton New-Style Quick-Dry Breathable Sports Competition Training
   Uniform”, ”brand”:”none”, ”category”:”Badminton Uniform”, ”price”:”88.00”, ”gmv”:”1625.0”,
   ”SID”:”<a_156><b_165><c_10><d_0>”},
 ”current query”:{”text”:”customized badminton rackets”, ”time”:”06/16 18:42:20”, ”location”:”Xining”}
Output:
¡think¿
The user is a male aged 31-40, indicating strong consumer capacity and a preference for high-end tech products, tending to choose powerful and reliable products…
The user exhibits a continuous interest in categories such as ”badminton equipment” and ”smartwatch accessories” based on their historical search behavior, especially with a high frequency of clicks on products from the ”YONEX” brand. This indicates a higher level of brand loyalty and a tendency to choose established brands or best-selling items…
The user shows strong purchasing power and tends to choose items with high cost-effectiveness, clear functions, and reputable brands within budget…
For the current query ‘customized badminton rackets’, the predicted item brand is YONEX, as the user has a clear preference for this brand’s products in the historical search…
Thus, the targeted product name should include keywords such as ‘new’, ‘professional’, ‘advanced’, and reflect high sales volume…
¡/think¿
<a_60><b_95><c_95><d_0>